What will be the wavelength of a ball of mass 0.1 kg moving with a velocity of 10 m s–1 ?

Grade:12

4 Answers

Gaurav
9 years ago
Hello Student
According to de Broglie equation
Wave length = h/mv = 6.255 x 10-34/(0.1x10) = 6.824 x 10-34 m
Ash Jain
6 years ago
h = 6.626 × 10^-34Now ,a/c to de- Broglie equation wavelength Wavelength = h/mvNow mv = 0.1×10=1So ,wavelength =h =6.626 ×10^-34 m
Sourav Singh
5 years ago
Acc. to the De Broglie equation,
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

According to de Broglie equation
Wave length = h/mv
= 6.255 x 10^-34/(0.1x10)
= 6.824 x 10^-34 m

Thanks and Regards